Privacy and Security Notice

Jefferson Lab
JLab Web Server 
Jefferson Lab Home Search Contact JLab

FEL Upgrade Project November 2002 Monthly Report

We had another very productive month.  Topping the list of major achievements includes getting rf in both cryomodules, one to full power with the other close behind.  We got the gun fully assembled with a cathode installed and ready for bakeout.  We completed the optical cavity assemblies and installed them in the tunnel.  We also put power in the wiggler and everything works.  Additional accomplishments with our installation and commissioning activities include the following:
(1) all four of the large dipole magnets (GW’s) for the first optical chicane were qualified by the magnetic measurement crew, found to meet our demanding specifications (1/1000), and installed in the FEL vault.  (2) the first reverse bend dipole (GQ) passed magnetic qualification;(3) we have identified several unique work-arounds for the late delivery of the recirculation hardware that allows our commissioning schedule to proceed (see WBS 9 write-up); with these work-arounds all beam transport for the first arc is now on site and being qualified for installation;  (4)  the High Voltage Power Supply tank for the electron gun was successfully pressure tested for operation.

An article was published in November in the prestigious journal “Nature” by Gwyn Williams et al that describes our successful measurement and analysis of the high power (20 watt), short pulse, terahertz radiation with IR Demo last year just before we shut down the device for FEL Upgrade installation.

We held the project Semiannual Review on Nov. 20-21 and appreciated the opportunity to show off the new machine to the review committee.

Nimel Theodore and Carlos Hernandez presented papers at the 49th AVS Symposium in Denver on Nov.4-8 on our ion implantation process used for preparing the gun electrodes. Nimel's contribution won second place in the student competition.

Significant effort was spent trying to resolve a severe cash flow crunch that has developed because of a significant curtailment of support from the State of Virginia, and, as of yet, non-receipt of JTO, AF, and most Navy funds for FY03.

Considerable effort was spent this month to analyze and document the FEL commissioning technical and cost status for the ONR Program Office.  Several scenarios were presented that prioritized FY03 FEL Upgrade activities with respect to potential FY03 funding levels.  These scenarios were reviewed at the Nov.20-21 FEL Upgrade Semiannual Review for the review committee’s advice and council.  We appreciate the time that was devoted to this analysis by Gil Graf at ONR and Wayne Skinner at our DOE Site Office.

The Semiannual Review on Nov. 20-21 was an exhausting event for both the presenters and the reviewers.  We appreciate the hard work delivered by all for this important event for the FEL Upgrade project.  We look forward to the committee’s report on their technical evaluation of our progress and we appreciate the support shown by both the reviewers and our DOE and DOD program managers for helping us to navigate the FY03 funding issues

George Neil and Steve Benson gave several presentations on the status of the FEL Upgrade program and scaling to higher power FEL systems at the Directed Energy Professional Society Symposium hosted by NAVSEA and the Naval Postgraduate School at Monterey, CA. on Nov. 12-14. FELs had very high visibility at the DEPS conference and in addition to George and Steve's papers, FEL Dept. members were listed as co-authors on three AES papers and two NPS papers.

Project Cost Performance:
The Phase 2 IR Demo FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002.  Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k.

We have met the planned cost at completion of IR Phase 2 effort.  All but 3% of the scheduled work (earned value) has been booked at the completion date.   Remaining items in the original project scope have not impeded the start-up of the FEL Upgrade commissioning which began in September. We are actively working with the vendors who are supplying the remaining hardware so that the delivery and integration of this hardware can be interleaved with the commissioning activities over the next 3 months and not impede the commissioning schedule.

The UV FEL project (Phase 1) was approved for $2,836k for a performance period from September 2001 to September 30, 2002.  Actual charges of $0k were accrued through November for a total accrued cost of $2,221k since project start.  Work performed for November was $157k for a total of $2,756k for the UV project to date, which results in 100% scheduled vs. 97% performed for the UV project.  At present the cost variance for the UV project (Phase 1) is $535k and the schedule variance is - $80k.  The current cost and schedule variances will clear when the last quarter of charges for the Aerospace subcontract are costed in December due to the 3 month late start on the subcontract.

Phase 2 of the UV FEL Project was approved for  $1,472k for a performance period from August 2, 2002 to June 30, 2003.  The start date was 4 months later than planned due to late arrival of AF funds.  Actual charges of $89k were accrued through November for a total accrued cost of $1,320k since project start.  Work scheduled for November was $272k for a total of $296k for the UV (Phase 2) project to date, work performed for November was $280k for a total of $1,353k for the UV (Phase 2) project to date, which results in 20% scheduled vs. 92% performed for the UV (Phase 2) project.  At present the cost variance for the UV (Phase 2) project is $33k and the schedule variance is $1,057k.  Because of the late arrival of funds significant expenditures had to made early in the first 2 months which have generated the large positive schedule variance.  The AFRL and DOE program offices agreed to FEL management's request in November to reprogram $414k held in the UV FEL Phase 2 account to support FEL commissioning tasks.  Tasks that were originally defined for an Aerospace subcontract under Phase 2 using these funds will be transferred to a Phase 3 Statement of Work when Phase 3 AF funds become available later in FY03.

WBS 3 (Beam Physics):
Results for various magnet measurements have been reviewed and the magnet’s performance certified.

Tune-up procedures are being simulated so as to provide final dump line layout specifications. An amusing result has emerged:  Conservation of energy limits the amount of energy you can recover.  (Corollary:  If the bunch length is longer than 360 degrees, you can't recover it very well...!)

WBS 4 (Injector):
The gun was pumped down, all ion pumps started and a satisfactory leak check was accomplished.  A small amount of contamination was discovered in the gun and injector line and a hot N2 purge was implemented to remove the bulk of it prior to the bake.

The two shielded bellows assemblies were pulled, solvent cleaned and vacuumed baked to remove oil contamination (at the 10E-10torr level) that was found during the initial RGA analysis.  The bellows were reinstalled, pumped down and the oil signature disappeared into the noise on the RGA.   A wafer was mounted and hydrogen cleaned.  The stalk with the wafer was mounted into the gun.  On pump down it was discovered the right angle valve had developed a leak across the seat at normal torque values and replaced. The gun is now fully assembled and under vacuum.  The external NF3 reservoir and line for wafer conditioning was installed.

The Conditioning Resistor assembly is completed. The HVPS Tank was also leak tested and passed. .  The extractor cart was removed from the ceramic stack and the corona rings and support rods were fit up.  The SF6 tank is being readied for fit up with the high voltage tank so we can mark the operating probe position and complete fabrication of the resistor.

WBS 6 (RF):
Buncher - The Buncher was warmed up without RF and the water loop operated.  All water fittings have been removed in preparation for the Gun and beam line bake.  One fitting interfered with the mounting bracket and had to be cut off.  A different style fitting is on order that will not interfere.

Zone 2 - SRF testing is complete with this zone.  It operates at its old Drive High values.  The new cavity tuning offset values are within +/- 5° of the old ones.

Zone 4 - SRF commissioning made good progress.  Nearly all cavities have been operated up to 10 MV/m.  The klystron cathode voltage was set too low (10.8 kV) and is being raised (11.6 kV) to allow the required RF power.  This was done, but several problems developed and are being addressed.

Quarter HVPS - The wiring for Unit #4 HVPS is complete and were tested for PSS Certification..  The klystron cart is wired, but still lacks some LCW fittings.  These fittings are on order.  The software appears to be working fine in all aspects.  The RF systems will be connected to the Quarter cavities for SRF testing.

WBS 8 (Instrumentation):
Trim magnet software and associated controls screens have been completed for the 1F to 3F region.  Epics communications for the 66 kW PS has been checked out successfully.  The PSS certification for the remaining large DC power supplies was completed.  DC power has been connected to the Wiggler and will be checked out during the lock up 11/8/02.  The new 5000 gallon bag for storage of sulfur hexafloride gas has been installed replacing the 2500 gallon bag used for the old gun HVPS.  Testing continues for the HVPS interlock/controller chassis. This will be installed next week.  The beam diagnostics for the 2G spectrometer line are being finalized.  A holder for the Si wafer (2G view screen) is being finalized with the shop.  A design for synchrotron light monitor attenuators is also being done with the shop.  We are building from stock Al parts rather than using a previous design that required extensive machining.

The diagnostics for the 2G spectrometer line are being assembled.  The design to hold a Si wafer (view screen) is being finalized with the shop.  Work progressed on the Wiggler viewers with fabrication, assembly and installation of the BV Interface boxes air will be connected early next week and the alignment will begin.

Drive Laser Pulse Controller (DLPC) - The VME version of the ConOptics Model 305 is built.  It picture of can be seen here:  This plug-in module will provide the function that the Model 305 chassis does and will be fully integrated into the ioc (iocfel10).  A link to a photo of the Model 305 chassis is here:  We are waiting on some rare RF connectors(SSMC, ~3 weeks) to finish the job.  The Model 305 is currently operational and will stay online until the new system is up and running.

Drive Laser Controls - Bringing all the functions of the Acromag card online continues.  This multifunction I/O card does three functions, namely:  discrete input and output (working), DAC control (working) & ADC input (in progress).  Additionally, work continues on the RF/picomotor controller board. Parts are on order and the CAD work in progress.

Identifying location and terminations of the signal and HV cables for the BLM system has begun in the 0F to 3F region.  System prints were submitted to EECAD for the magnet communication upgrade as well as the new magnet configuration in the FEL.  We continue to work with EECAD to clear up the backlogged drawings in their queue.

Drive Laser Controls - The acromag card is fully implemented into iocfel10 and all functions (Digital-I/O, ADC & DAC channels) are fully checked-out.

Drive Laser Pulse Controller (DLPC) Upgrade - The SSMC-RF connectors for the CO305 VME Module came in sooner than expected and so the board is being finished for testing. Also, as part of the DLPC upgrade, a full system print of the drive laser as a sub-system is being produced.  This will document everything in the drive laser clean room from an as-built/controls perspective.  With the basic connectivity of the drive laser's RF controls documented including the custom modifications to the Modelocker chassis, we are in a good position to lock-in the design of the automatic switch to internal RF as a part of the external RF interlock upgrade.  This control function prevents the possibility of a "Q-Switch" and cooking the Second Harmonic Generation (SHG) crystal.  The modified Modelocker chassis with the manual RF switch is this weeks picture link at:

Beam Loss Monitors (BLMs) - The F0151 module has been pulled from iocfel8 and given to EES as a the functional prototype to aid in the assembly of the rest of the boards.  We will get it back as they finish new boards one at a time versus assembling the lot in parallel.

The new beam viewer design is complete and being manufactured (THANKS Casey!). This design is a 1"x4" silicon paddle fixed to a 6" conflat blank.  The paddle has registration marks at 5mm with a major mark on center.  There are two of these being built one for the strait ahead spectrometer and the second for the GW beam diverter downstream of the wiggler.  All parts are on order for the Synchrotron Light Monitor (SLM) attenuators.  This is a very simple design based on 2"x4" box aluminum beam.  Assembly of the prototype should take place next week.

The old OPCHIC DC power supply has been reinstalled for powering the DW & DY (work around).  It was moved into place and had its AC power connected.  The FEL I&C team has wired this into the PSS to enable certification this month during the scheduled PSS certification of the 1/4CM HPA.  This is the LAST PSS connected power supply. The software group has begun work on the software and screens to support the old OPCHIC power supply.  DC power to the 0F region corrector magnets is near completion.  The new GW does not require a PSS connection since it has the leads covered and is only a 5 kW supply.
The optical cavity wiring (controls) is progressing; much of this activity is done after 1700 and must be coordinated with the (nearly) nightly lockups for RF commissioning.  The control screen for the 4 channel BPMs are being worked.
Beamviewers for the wiggler have been installed and checkout is in progress.  The 25 pin MPS breakout modules have been received from the vendor and are being populated in house. Beamviewer MPS cables were pulled into FL02B10 and termination is in progress.  Work progressed on the BLM high voltage cable terminations.

Schematic capture for the HV Interlock digital I/O board have been received from EECAD and under review.  Work has progressed on the system print for the HV interlock.  The software for the Machine Protection System (MPS) is progressing very well.  There are some screens done with the new cards all reading back.  Many of the input cables are complete and ready to connect, this will be done as priorities allow.

Magnet shunt cable for the 1st, 2nd arc strings, injection dipole, extraction dipole and optical chicane have been pulled into their respective zones.  Termination to the shunt chassis is in progress.  Work has begun on installing the 200 amp DC power leads for the 1st and 2nd arc magnet strings.

BLM signal and HV cables have been terminated in the MPS rack. Reconfiguration of the MPS rack is in progress.  Fabrication of the HV Interlock chassis is complete and functional testing has begun.

Re-termination of the 4 channel BPM drops (1F01 thru 3F11) have been completed. The 25 pin MPS breakout modules were populated.  The beamviewer MPS interlock cables were routed between zones and terminated.  Cable termination's for the high reflector optical cavity has begun in the vault.  Work has begun on installing the magnet shunt cable for the 1st, 2nd Arc strings and optical chicane.  The magnet thermal interlock chassis has been installed in Zone 5 and awaiting communications hook up for check out.

Schematic capture for the HV Interlock analog I/O board has been received from Physics and is under revision.  EECAD has begun front panel fabrication drawings for the 15 and 25 pin MPS breakout modules.  The 1st revision of the BLM system wiring diagram was submitted to EECAD for mark ups.  System wiring mark ups for the Magnet Com. upgrade and Magnet strings have been received and under review.

The testing of the optical cavity position sensors (LVDTs) has begun. A bad channel was found on one of the VME cards, it is being sent out for repair (we have a spare). Cabling is complete for the High Reflector end; the motors and encoder (only on 'lift') will be tested as they become available.

WBS 9 (Transport):
Work Arounds - Work arounds are now implemented to compensate for the recirculation magnets that are late while maintaining the commissioning schedule. Dave Douglas has blessed all these efforts as worthy of pursuit.

1. To commission the linac without the first arc complete in December, we need a dipole to send the beam off at an angle at the end of the linac to phase the cavities.  We installed an IR Demo Optical Chicane Dipole and its chamber just after the GX position with an innovative silicon wafer beam viewer developed by Kevin Jordan.  It will be powered by the spare 66 kW supply in series with the DY below. Control was simplified by powering both provisional magnets on the same supply because machine commissioning uses only one of these magnets at any one time.

2. For first light commissioning using pulsed beam, mid January, use an IR Demo 180° Dipole (DY) to turn the first arc (along with the GQ and the GX dipoles that will be measured by then). Fortunately DY is rated for 80 MeV/c (the commissioning momentum) and its 2-inch aperture is adequate for the requirements of the first arc.  The DY only needs new flanges on its chamber to adapt to the GQ-GX chambers.

We will extract pulsed electron beam using a GW dipole three quadrupoles past the optical klystron so we don’t hit the mirror with the electron beam.  At that position, we can use an IR Demo optical chicane chamber for the 20° bend-out and we can easily stack SEG (steel shielding) blocks to absorb the beam.  The GW dipole will be one of the dipoles intended for the UV Line.  We will power it with the already installed power supply for the extraction chicane, (which is not powered during this phase of the effort).

3. For recirculation and high power lasing, starting near the end of February, install the 1st 180° dipole and GQs & GXs in the second arc with the IR Demo’s DY remaining in the first Arc and remove the GW above.

4. Since the sextupoles are designed but not procured, work around using IR Demo SC sextupoles. We are expanding them with shim blocks and pole shifts to fit over larger chamber and they are just strong enough for the commissioning energy of 80 MeV if we double their current and cool

Optical Chicane Dipoles (GW)
• Continued magnetic success. Magnetic tests on the stepper stand qualified all four GWs for the IR Machine.  They meet David Douglas’ specifications for core field flatness in the transverse (± 1 part in 10,000), field integral over the good field region (± 1 part in 10,000) and exceeded the spec by a factor of two for core field and field integral within the magnet family (within a part in 1000) at the nominal machine energy of 145 MeV. All four were released and installed this month
• Additional tests verified the levels of flatness on the “Dipole” Stand where a loop-coil tests field integral flatness.  The Group is using these magnets to calibrate the stand for use on the GQ-GX styles, coming next.
• Kudos to Magnet Enterprises International and the collection of machining and grinding vendors in the Oakland industrial area for making these magnets so beautifully, especially gluing the Purcell gap material down to the pole tips to the required precision.
• Kudos also to the Magnetic Measurement Group and their help from the JLab Machine Shop for running a 16 hr. schedule and measuring the 4 GWs all in one week, working like clock work.
• At MEI, one of the 6 UV GWs is complete, two have the Purcell gap pole tips completely glued down and qualified and one more has its top shims being glued.  They plan to ship in a week and one half.
Injector Dipoles (DU/DV)
• No new developments
Arc 180 Degree Dipoles (GY)
• Wang NMR is just about ready to pot the first coil and continues to set up for gluing of Purcell gap shims. They do not want to spoil these operations.
• We visited them to witness these critical operations.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
• Two GQs had their LCW hoses applied by the machinist who is supporting the magnet measurement effort. They are being prepared for measurement at the start of next week. They look good.
• Two GXs will follow.
• Wang continues to vacuum impregnate the new series of coils for the remaining two GX magnets and is converting the mold for the GQ coils to this method.

Trim Quad (QT)
• Trim quads were sent to the FEL for installation and trial fit up.  Four will have to be retrofitted with coils to add the GC corrector function to them.
Sextupole (SF)
• Master Machine completed one “Commissioning Sextupole” for trial installation in the enclosure. The remaining 7 are nearly complete.
Octupole (OT)
• We continue working on the procurement package to get these 2 magnets made in pieces by local shops. We have the coils already.
Beam Line and Vacuum
• Chambers for the arcs continue in production status.  We received four chambers this month.  The first arc’s chambers are nearly complete.  Master Machine notified us that they would not make their end of November deadline to complete the task for the second arc. The completion will be first week of December.  This does not trail the magnet availability schedule, so we should be all right.
• The vacuum pipes between girders leading to the wiggler region were installed.
• The magnet measurement crew stopped installation of the correctors to concentrate on GWs.
• LCW piping fabrication is at 95%.
• The Dump line stand positions (with a variable parameter of the dump face) were issued to the alignment group to start installation. David Douglas is ruminating on the best compromise position for the face that allows the best control of betatron and disbursive beam qualities.
• Layouts of the workarounds above are nearly complete.

WBS 10 (Wiggler):
Ran the wiggler with its power supplies.  Power supplies, PSS, etc., all worked great.

WBS 11 (Optics):
This month’s activities culminated with our transporting the optical cavity vacuum vessels to the FEL vault.  The support girders with ultraviewers attached, were installed next to their vacuum vessels and aligned by the Survey and Alignment team.  A number of incremental milestones, were accomplished to do this, the major ones are listed below:

Working with Survey and Alignment, the internal optical components for both the HR and OC assemblies were fiducialized in the vacuum vessels.  We then translated the mirrors for both assemblies, and checked the repeatability.  The data reflected the repeatability of the portable CMM, ~ 100 microns.  That is our specification.  There is evidence that the performance is higher.  The HR vacuum vessel, with internal hardware for vertical motion, was leak checked, and checked for oil contamination.  It passed both tests.  Clean enclosures were established for both the HR and OC assemblies before they were moved.

During the above, we also completed inspection of the first light outcouplers.  One failed (visible flaws within 80% CA) but the other passed.  We also began installing one of the two alignment HeNe lasers we'll use for our alignment activities.

The custom optical breadboards for the OCMMS optics arrived.  Drawings for the insertable mirror used to route high power beam to the diagnostic beam dump are being checked.  The vacuum vessels for the high power optical transport are in Procurement but purchase is on hold pending additional funds (this item was descoped from the construction project).

Due to a ~ 2 week delay in Customs, we anticipate receiving the rest of the UHV linear actuators to remotely position the cavity mirrors on Dec 2.  This order also includes the actuators for the collimator and most of the optical transport to the optics control room.  We received one of the 800 nm HR optics we'll use to set the cavity length.

We signed off on drawings for the laser safety shutter and the insertable mirror we'll use for wiggler spontaneous emission.  At the same time we edited drawings for the insertable mirror used to route high power beam to the diagnostic beam dump.  We are in the process of planning the installation of laser diagnostics in the FEL vault.  We received quotes on an absorptive coating for the high power diagnostic beam dump.  This process looks promising, but there is some uncertainty on power-handling capability, so we have requested a sample from the vendor.  We'll test it in the Mirror Test Stand.

Aerospace/ Jefferson Lab Subcontract
(for the Laser Microengineering Experimental Station)
We continued to procure equipment for the Engineering Model subsystems (optical table, XYZ motion control, ps-laser, imaging system and custom-design pockels cells).  Purchase orders were placed for the optical table with active stabilization, the ps-laser that will mimic the single pulse energy of the UV FEL, XYZ motion control system and a pair of Pockels cells for the pulse picking and intensity control.  The goal is to have all purchase order signed by the end of Contract date (December, 11, 2002).

A presentation was given at the JLAB FEL PI meeting.  The presentation included details of the Engineering Model design, the impact of the expected Phase Ib (12-24 months segment) funding shortfall and a rough draft of the modified statement of work (SOW) for Phase Ib.  An additional three copies of the preliminary design document (discussed in the August Report) was assembled and delivered to JLAB staff.   The presentation also stressed the urgency of getting the FY03 funds released to Aerospace by March time frame.  Additional delay in funding will result in loss of manpower availability and the incurring of additional cost to outsource the tasks in the Phase Ib SOW.

Other Related Projects

IR Upgrade Commissioning (ONR)
This project was funded in FY02 with $1,000k, with a period of performance from July 1, 2002 to March 30, 2003. Additional funding of $1,000k is anticipated in FY03 along with a requested re-programming of unexpended FY02 funds and requested FY03 funds for the one micron wiggler project, which will be deferred until completion of FEL commissioning.  In this reporting period, $1,066k has been expended for a total expenditure of $1,708k.  Commitments total $69.6k.
See Highlights section and WBS 4,6,8,9,10 and 11 reports above.

Cryomodule 3 (ONR)
This project was funded in FY02 with $400k.  Additional funding of $350k is anticipated in FY03.  In this reporting period, $189k has been expended for a total expenditure of $422k.  Commitments total $2k.
See WBS 5 report above.

One micron wiggler project (ONR)
This project was funded in FY02 with $600k.  Completion funding of $600k is anticipated in FY03.  In this reporting period, $87k has been expended for a total expenditure of $209k.  Commitments total $10.3k.
Prepared presentation for Navy closeout review on this subject and deferred further activities pending our recommendation to defer all further work on this task until the FEL Upgrade is commissioned. This is a good stopping point because the dominant cost at this point is the capital expenditure for the one micron wiggler.

JTO Cryomirror Project
During this reporting period expenditures of $113.5k were incurred including bookings of previous commitments.  Commitments now total $41.2k.  Total expenditures to date are $261k.  The program is on track financially and is expected to deliver all elements of the SOW within projected cost.  It is expected that this budget will be fully expended in December with Phase 1 elements complete.  We are prepared to continue on schedule with Phase 2 of the program.
We proceeded with engineering tests of the proposed system.  The efforts performed essentially completed the SOW for this program with the exception of the final report.
We performed measurements of the distortion of sapphire at cryogenic temperatures under two heat loads at three temperatures corresponding to levels at which a 100 kW and a MW class FEL system would operate.  The measurements were entirely successful and confirmed predictions of low distortion under these thermal loads.  Interferograms were presented which illustrate the improvement as the temperature is lowered to 100K.  This actually represents a worst case scenario since the planned operating temperature is below 70K where the material parameters are further improved.

JTO Drive Laser Project
During this reporting period expenditures of $19.6 k were incurred.  Commitments total $0k.  Total expenditures to date are $43k.  The program is on track financially and is expected to deliver all elements of the SOW within projected cost. It is expected that this budget will be fully expended in December with Phase 1 elements complete.  We are prepared to continue on schedule with Phase 2 of the program.
We are in the process of proceeding on the major laser hardware purchase.
We sent out the purchase order for the master oscillator system.